# Geodesic Synthetic Control Methods for Random Objects and Functional Data

### v1

@@ Abstract.
Key words and phrases:
@@
causal inference, distributions, Fréchet mean, geodesic metric space, networks, synthetic difference-in-differences
MSC2020 subject classifications: 62D20, 62R20
The work of D. K. was partially supported by JSPS KAKENHI Grant Number 23K12456 and the work of H.G.M. was partially supported by NSF grant DMS-2310450
1. Introduction
@@
Since the seminal work by Abadie and Gardeazabal, (2003) and Abadie et al., (2010), synthetic control methods (SCMs) have been widely applied to evaluate causal effects of policy changes in settings with observational panel or longitudinal data. SCMs are designed for settings where some units are subject to a policy intervention while others are not, and their outcomes are observed before and after the intervention. In such settings, synthetic control units constructed utilizing some optimal weighting of control units are employed to control for unobserved trends in the outcome over time that are unrelated to the policy effect. The SCM has become popular in many areas and has been applied in various comparative case studies of policy interventions (see Abadie, (2021) for a survey).
@@
Following the above pioneering work, the SCM has been extended in various directions, including an SCM that incorporates a penalty term for pairwise matching discrepancies between the treated units and the control units (Abadie and L’Hour,, 2021), a SCM for high-dimensional settings (Robbins et al.,, 2017), a connection with matrix completion methods (Amjad et al.,, 2018) and conformal inference and robust $t$-testing for SCMs (Chernozhukov et al., 2021a, ; Chernozhukov et al., 2021b, ). A modification of the synthetic control estimator by incorporating a bias correction method for inexact matching was proposed by Ben-Michael et al., (2021) and for a staggered adoption setting by Ben-Michael et al., (2022), while Arkhangelsky et al., (2021) introduced a synthetic difference-in-differences (SDID) estimator that combines attractive features of the SCM and difference-in-differences (DID).
@@
However, all of these developments primarily focused on Euclidean outcomes. In modern data analysis, one frequently encounters outcomes that may be inherently non-Euclidean, including distributional data (Petersen and Müller,, 2016; Panaretos and Zemel,, 2020; Petersen et al.,, 2022), compositional data (Scealy and Welsh,, 2011), manifold-valued data (Patrangenaru and Ellingson,, 2015), networks (Severn et al.,, 2022; Zhou and Müller,, 2022) and functional data (Hsing and Eubank,, 2015; Wang et al.,, 2016). Such data reside naturally in metric spaces where one often imposes additional mild properties such as geodesicity, which is satisfied for nearly all outcomes of interest, including those mentioned above under appropriate metrics; common choices are the Wasserstein metric for distributional spaces, the Riemannian metric on spheres for compositional data and on more general manifolds for manifold-valued data, the Frobenius or power metric when representing networks through their graph Laplacians, and the $L^{2}$ metric for the case of functional data. The Frobenius and power metrics, among other metrics, are also commonly used for the space of symmetric positive-definite matrices, which may also appear as outcomes, for example when correlation structures between random vectors are outcomes of interest. Another type of potential interest is the tree spaces with the BHV metric (Billera et al.,, 2001). All of these are geodesic metric spaces, requiring specialized approaches for causal inference. The challenge here is that these spaces, with the exception of the Hilbert space where functional data are situated, are inherently nonlinear, i.e., one does not have available linear operations such as addition, subtraction and scalar multiplication on which most established methods rely. General approaches to address these challenges for data analysis in metric spaces are discussed in Dubey et al., (2024).
@@
Addressing causal inference for outcomes in geodesic spaces is a recent but rapidly evolving area. Recent studies include estimators for average treatment effects (Lin et al.,, 2023), DID (Torous et al.,, 2024), SCMs (Gunsilius,, 2023; Gunsilius et al.,, 2024), and regression discontinuity designs (Van Dijcke,, 2025); all of these papers focus on the special case of distributional outcomes. Recent contributions addressing more general outcomes that can be viewed as elements of a general geodesic metric space include Kurisu et al., (2024) and Zhou et al., (2025), where in the latter a geodesic DID (GDID) approach was developed.
@@
In this paper, we introduce a geodesic synthetic control method (GSC) designed for outcomes residing in geodesic metric spaces such as those mentioned above. The GSC framework thus accommodates a broad array of outcomes residing in geodesic metric spaces. For the special case of distributional outcomes, GSC encompasses the distributional synthetic control framework of Gunsilius, (2023) when using the Wasserstein metric. However, GSC also permits the use of alternative metrics in distribution space, as long as the metric space is geodesic. This includes the Fisher-Rao metric and many other distribution metrics of interest, thereby broadening the scope of analysis even for the special case of distributional outcomes.
@@
We introduce a causal model suitable for the GSC and establish theoretical identification results for the causal effects of interest. Specific examples of causal models satisfying our assumptions in various common geodesic metric spaces illustrate the practical applicability of the proposed methodology. Additionally, we develop an augmented GSC, motivated by the augmented SCM introduced by Ben-Michael et al., (2021) for Euclidean outcomes. This extension is non-trivial due to the absence of linear algebraic structures in geodesic metric spaces; it necessitates the use of geodesic transport maps to construct our estimators.
@@
We further propose a geodesic synthetic difference-in-differences (GSDID) estimator, extending the SDID methodology introduced by Arkhangelsky et al., (2021) for Euclidean outcomes. The GSDID estimator combines desirable properties of both the proposed GSC and the GDID developed by Zhou et al., (2025), which exhibits a double robustness property: even if the weighting schemes from either GSC or GDID individually fail to eliminate bias, their combination within the GSDID framework provides additional robustness. Moreover, the GSDID framework subsumes both the GSC and GDID as special cases.
@@
The paper is organized as follows. Section 2 provides a concise review of metric geometry and introduces several geodesic metric spaces relevant for simulations and empirical analyses. Section 3 presents the proposed GSC framework, including identification results and an extension to augmented GSC. Implementation details and simulation studies for network and symmetric positive-definite matrix outcomes are described in Section 4. Section 5 illustrates the practical applicability of GSC through two empirical examples: analyzing changes in employment composition following the 2011 Great East Japan Earthquake, and assessing the impact of abortion liberalization on fertility patterns in East Germany. Section 6 extends the methodology to GSDID, presenting identification results along with an illustrative example investigating the health consequences of the Soviet Union’s collapse through age-at-death distributions. Finally, Section 7 concludes the paper. Proofs of theoretical results are provided in the Appendix.
2. Preliminaries on Metric Geometry
@@
where the supremum is taken over all possible finite partitions $10=t_{0}\leq\cdots\leq t_{I}=$ 1. A curve $\gamma$ is called a geodesic if it satisfies $d(\gamma(s),\gamma(t))=|t-s|d(\gamma(0),\gamma(1))$ for all $s,t\in[0,1]$. The metric space $\mathcal{M}$ is called a geodesic space if, for any two points $\alpha,\beta\in\mathcal{M}$, there exists a geodesic connecting them, denoted $\gamma_{\alpha,\beta}$. If this geodesic is unique for every pair of points in $\mathcal{M}$, the space is referred to as a unique geodesic space (Bridson and Haefliger,, 1999). A basic assumption for our approach is that the outcomes reside in such a space.
@@
Example 2.1 (Networks).
Consider the space of simple, undirected, weighted networks with $m$ nodes and bounded edge weights. Each network can be uniquely represented by its graph Laplacian matrix. The space of graph Laplacians $\mathcal{L}$ serves as a natural framework for characterizing network structures (Kolaczyk and Csárdi,, 2014; Severn et al.,, 2022; Zhou and Müller,, 2022). When endowed with the Frobenius metric, defined as $d_{F}(L_{1},L_{2})=\|L_{1}-L_{2}\|_{F}$ for $L_{1},L_{2}\in\mathcal{L}$, the space $\mathcal{L}$ becomes a unique geodesic space. In this setting, the geodesic between $L_{1}$ and $L_{2}$ is given by the linear interpolation $\gamma_{L_{1},L_{2}}(t)=(1-t)L_{1}+tL_{2}$ for $t\in[0,1]$.
Example 2.2 (Compositional Data).
@@
Adopting the square-root transformation provides a map from $\Delta^{d-1}$ to the positive orthant of the unit sphere $\mathcal{S}_{+}^{d-1}$ (Scealy and Welsh,, 2011), where the induced geodesic distance on $\mathcal{S}_{+}^{d-1}$ is the arc-length distance
@@
[…]
@@
Example 2.3 (Symmetric Positive-Definite Matrices).
The space of $m\times m$ symmetric positive-definite (SPD) matrices, denoted by $\mathrm{Sym}^{+}_{m}$, plays a crucial role in various fields such as computer vision (Cherian and Sra,, 2016), signal processing (Arnaudon et al.,, 2013) and neuroimaging (Dryden et al.,, 2009). Various metrics have been introduced to endow $\mathrm{Sym}^{+}_{m}$ with a meaningful geometric structure, depending on the application. The Frobenius metric $d_{F}$ induces unique geodesics that are simple linear interpolations, $\gamma_{A,B}(t)=(1-t)A+tB$ for $A,B\in\mathrm{Sym}^{+}_{m}$ and $t\in[0,1]$. The space $\mathrm{Sym}^{+}_{m}$ also forms a unique geodesic space under various other metrics such as the Log-Euclidean metric (Arsigny et al.,, 2007), the power metric family (Dryden et al.,, 2009) or the Log-Cholesky metric (Lin,, 2019). For instance, when adopting the log-Euclidean metric, the unique geodesic connecting $A,B\in\mathrm{Sym}^{+}_{m}$ is given by
@@
Example 2.4 (One-Dimensional Probability Distributions).
@@
[…]
where $F_{\mu}^{-1}$ and $F_{\nu}^{-1}$ are the quantile functions of the probability measures $\mu$ and $\nu$, respectively. The Wasserstein space $\mathcal{W}$ is a complete, separable and unique geodesic space (Ambrosio et al.,, 2008), with geodesics given by McCann’s interpolant (McCann,, 1997)
@@
Example 2.5 (Functional Data).
In functional data analysis, one considers data consisting of functions that are considered to be the realizations of a square integrable stochastic process, usually assumed to be defined on a finite interval. Such data are common in the social, physical and life sciences (Ramsay and Silverman,, 2005; Hsing and Eubank,, 2015; Wang et al.,, 2016). Typically, functional data are assumed to be situated in the Hilbert space $L^{2}(\mathcal{T})$ of square integrable functions, where $\mathcal{T}$ is a compact interval, most commonly a time domain. In contrast to the other metric spaces we consider in this paper, this space is a vector space and thus permits algebraic operations. It is equipped with the standard $L^{2}$ metric
[…]
@@
3. Geodesic Synthetic Control Methods
3.1. Construction of the estimator
@@
Consider a standard panel data setting for the synthetic control method (SCM), where outcomes are observed for $J+1$ units over $T$ time periods, indexed by $j=1,\dots,J+1$ and $t=1,\dots,T$, respectively. In what follows, we fix both $J$ and $T$. For Euclidean outcomes $Y_{j,t}\in\mathbb{R}^{q}$, the conventional SCM is as follows. Assume that for a time $T_{0}>1$ only the unit indexed by $j=1$ experienced the intervention at times $t$ with $T_{0}<t\leq T$. We denote the potential outcomes of unit $j$ at time $t$ with and without intervention by $Y_{j,t}(I)$ and $Y_{j,t}(N)$, respectively. We observe $Y_{j,t}=Y_{j,t}(I)$ only for $j=1$ and $t=T_{0}+1,\ldots,T$. For all other combinations of $j$ and $t$ we observe $Y_{j,t}=Y_{j,t}(N)$. The causal effect of interest is $\tau_{t}=Y_{1,t}(I)-Y_{1,t}(N)=Y_{1,t}-Y_{1,t}(N)$ for $t=T_{0}+1,\ldots,T$. The SCM is a method to predict $Y_{1,t}(N)$ for $t=T_{0}+1,\ldots,T$ using the control units $\{Y_{j,t}=Y_{j,t}(N)\}_{2\leq j\leq J+1,1\leq t\leq T}$.
In the SCM, the predicted value for $Y_{1,t}(N)$ is constructed as
$Y_{1,t}^{(\mathrm{SC})}:=\sum_{j=2}^{J+1}\overline{w}_{j}Y_{j,t}=\mathop{\rm arg~{}min}\limits_{y\in\mathbb{R}^{q}}\sum_{j=2}^{J+1}\overline{w}_{j}\|y-Y_{j,t}\|^{2},\quad t=T_{0}+1,\dots,T,$
where the weights are given by
@@
$\overline{\bm{w}}=(\overline{w}_{2},\dots,\overline{w}_{J+1})^{\prime}\in\mathop{\rm arg~{}min}\limits_{\bm{w}\in\Delta^{J-1}}\frac{1}{T_{0}}\sum_{t=1}^{T_{0}}\left\|Y_{1,t}-Y_{2:J+1,t}^{(\bm{w})}\right\|^{2},$
$Y_{2:J+1,t}^{(\bm{w})}=\sum_{j=2}^{J+1}w_{j}Y_{j,t}=\mathop{\rm arg~{}min}\limits_{y\in\mathbb{R}^{q}}\sum_{j=2}^{J+1}w_{j}\|y-Y_{j,t}\|^{2},$
$\Delta^{J-1}=\{\bm{w}=(w_{2},\dots,w_{J+1})^{\prime}\in\mathbb{R}^{J}:w_{j}\geq 0,\sum_{j=2}^{J+1}w_{j}=1\},$
and thus are based only on the observations made at times $t$ with $1\leq t\leq T_{0}$. This motivates to obtain the estimate of the causal effect $\tau_{t}$ as $\tau_{t}^{(\mathrm{SC})}=Y_{1,t}-Y_{1,t}^{(\mathrm{SC})}$. The estimators $\{\tau_{t}^{(\mathrm{SC})}\}_{t=T_{0}+1,\dots,T}$ are referred to synthetic control estimators.
To extend the SCM for outcomes residing in a unique geodesic space $(\mathcal{M},d)$ one needs to substitute the algebraic operation of taking a difference that is crucial for the definition of the estimator $\tau_{t}^{(\mathrm{SC})}$, since one cannot take sums or differences of random objects in general metric spaces. Here, the assumption that the underlying space is geodesic becomes essential. In Euclidean spaces, the difference $\beta-\alpha$ can be equivalently interpreted as movement along the geodesic (straight line) connecting $\alpha$ to $\beta$. This insight motivates the generalization of differences to geodesics connecting corresponding points (Kurisu et al.,, 2024). Accordingly, we introduce the causal effect of the intervention in a unique geodesic space as the geodesic from $Y_{1,t}(N)$ to $Y_{1,t}(I)$, i.e.,
$\tau_{t}=\gamma_{Y_{1,t}(N),Y_{1,t}(I)},\quad t=T_{0}+1,\dots,T.$ (3.1)
To provide additional motivation why this makes sense, in the Euclidean (vector) case the difference $Y_{1,t}(I)-Y_{1,t}(N)$ utilizes both the shortest distance and directional information when moving from $Y_{1,t}(N)$ to $Y_{1,t}(I)$. Since a geodesic connecting points $\alpha$ and $\beta$ in a geodesic metric space $(\mathcal{M},d)$ embodies both the shortest distance and directional information from $\alpha$ to $\beta$ with respect to the metric $d$, $\tau_{t}$ emerges as a natural extension of quantifying treatment effects in geodesic metric spaces.
To estimate the causal effect $\tau_{t}$ in (3.1), we adopt a natural extension of the construction of synthetic weights in the Euclidean case and then obtain the predicted value for $Y_{1,t}(N)$ as $Y_{1,t}^{(\mathrm{GSC})}=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}\overline{w}_{j}d^{2}(\nu,Y_{j,t})$, where
$\overline{\bm{w}}\in\mathop{\rm arg~{}min}\limits_{\bm{w}\in\Delta^{J-1}}\frac{1}{T_{0}}\sum_{t=1}^{T_{0}}d^{2}(Y_{1,t},Y_{2:J+1,t}^{(\bm{w})}),$ (3.2)
$Y_{2:J+1,t}^{(\bm{w})}=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,Y_{j,t}).$
The resulting geodesic synthetic control (GSC) estimator in unique geodesic spaces is

### v2

@@ Abstract
11footnotetext: The first two authors contributed equally to this work and are listed alphabetically.22footnotetext: Corresponding author: hgmueller@ucdavis.edu.33footnotetext: D.K. was supported in part by JSPS KAKENHI Grant Number 23K12456 and 25K00624. H.G.M. was partially supported by NSF grant DMS-2310450.
@@
Keywords: causal inference, Fréchet mean, geodesic metric space, program evaluation, synthetic difference-in-differences
1 Introduction
@@
Since the seminal work by Abadie and Gardeazabal (2003) and Abadie et al. (2010), synthetic control methods (SCMs) have been widely applied to evaluate causal effects of policy changes in settings with observational panel or longitudinal data. SCMs are designed for settings where some units are subject to a policy intervention while others are not, and their outcomes are observed before and after the intervention. In such settings, synthetic control units constructed utilizing some optimal weighting of control units are employed to control for unobserved trends in the outcome over time that are unrelated to the policy effect. The SCM has become popular in many areas and has been applied in various comparative case studies of policy interventions (see Abadie (2021) for a survey).
@@
Following the above pioneering work, the SCM has been extended in various directions, including an SCM that incorporates a penalty term for pairwise matching discrepancies between the treated units and the control units (Abadie and L’Hour, 2021), a SCM for high-dimensional settings (Robbins et al., 2017), a connection with matrix completion methods (Amjad et al., 2018) and conformal inference and robust $t$-testing for SCMs (Chernozhukov et al., 2021a, b). A modification of the synthetic control estimator by incorporating a bias correction method for inexact matching was proposed by Ben-Michael et al. (2021) and for a staggered adoption setting by Ben-Michael et al. (2022), while Arkhangelsky et al. (2021) introduced a synthetic difference-in-differences (SDID) estimator that combines attractive features of the SCM and difference-in-differences (DID).
@@
However, all of these developments primarily focused on Euclidean outcomes. In modern data analysis, one frequently encounters outcomes that may be inherently non-Euclidean, including distributional data (Petersen and Müller, 2016; Panaretos and Zemel, 2020; Petersen et al., 2022), compositional data (Scealy and Welsh, 2011), manifold-valued data (Patrangenaru and Ellingson, 2015), networks (Severn et al., 2022; Zhou and Müller, 2022) and functional data (Hsing and Eubank, 2015; Wang et al., 2016). Such data reside naturally in metric spaces where one often imposes additional mild properties such as geodesicity, which is satisfied for nearly all outcomes of interest, including those mentioned above under appropriate metrics; common choices are the Wasserstein metric for distributional spaces, the Riemannian metric on spheres for compositional data and on more general manifolds for manifold-valued data, the Frobenius or power metric when representing networks through their graph Laplacians, and the $L^{2}$ metric for the case of functional data. The Frobenius and power metrics, among other metrics, are also commonly used for the space of symmetric positive-definite matrices, which may also appear as outcomes, for example when correlation structures between random vectors are outcomes of interest. Another type of potential interest is the tree spaces with the BHV metric (Billera et al., 2001). All of these are geodesic metric spaces, requiring specialized approaches for causal inference. The challenge here is that these spaces, with the exception of the Hilbert space where functional data are situated, are inherently nonlinear, i.e., one does not have available linear operations such as addition, subtraction and scalar multiplication on which most established methods rely. General approaches to address these challenges for data analysis in metric spaces are discussed in Dubey et al. (2024).
@@
Addressing causal inference for outcomes in geodesic spaces is a recent but rapidly evolving area. Recent studies include estimators for average treatment effects (Lin et al., 2023), DID (Torous et al., 2024), SCMs (Gunsilius, 2023; Gunsilius et al., 2024), and regression discontinuity designs (Van Dijcke, 2025); all of these papers focus on the special case of distributional outcomes. Recent contributions addressing more general outcomes that can be viewed as elements of a general geodesic metric space include Kurisu et al. (2024) and Zhou et al. (2025), where in the latter a geodesic DID (GDID) approach was developed.
@@
In this paper, we introduce a geodesic synthetic control method (GSC) designed for outcomes residing in geodesic metric spaces such as those mentioned above. The GSC framework thus accommodates a broad array of outcomes residing in geodesic metric spaces. For the special case of distributional outcomes, GSC encompasses the distributional synthetic control framework of Gunsilius (2023) when using the Wasserstein metric. However, GSC also permits the use of alternative metrics in distribution space, as long as the metric space is geodesic. This includes the Fisher-Rao metric and many other distribution metrics of interest, thereby broadening the scope of analysis even for the special case of distributional outcomes.
@@
We introduce a causal model suitable for the GSC and establish theoretical identification results for the causal effects of interest. Specific examples of causal models satisfying our assumptions in various common geodesic metric spaces illustrate the practical applicability of the proposed methodology. Additionally, we develop an augmented GSC, motivated by the augmented SCM introduced by Ben-Michael et al. (2021) for Euclidean outcomes. This extension is non-trivial due to the absence of linear algebraic structures in geodesic metric spaces; it necessitates the use of geodesic transport maps to construct our estimators.
@@
We further propose a geodesic synthetic difference-in-differences (GSDID) estimator, extending the SDID methodology introduced by Arkhangelsky et al. (2021) for Euclidean outcomes. The GSDID estimator combines desirable properties of both the proposed GSC and the GDID developed by Zhou et al. (2025), which exhibits a double robustness property: even if the weighting schemes from either GSC or GDID individually fail to eliminate bias, their combination within the GSDID framework provides additional robustness. Moreover, the GSDID framework subsumes both the GSC and GDID as special cases.
@@
The paper is organized as follows. Section 2 provides a concise review of metric geometry and introduces several geodesic metric spaces relevant for simulations and empirical analyses. Section 3 presents the proposed GSC framework, including identification results and extensions to augmented GSC and GSDID. Implementation details and simulation studies for network and symmetric positive-definite matrix outcomes are described in Section 4. Section 5 illustrates the practical applicability of GSC through two empirical examples: analyzing changes in employment composition following the 2011 Great East Japan Earthquake, and assessing the impact of abortion liberalization on fertility patterns in East Germany. Finally, Section 6 concludes the paper. Proofs of the main theoretical results and additional materials are provided in the Supplementary Material.
2 Preliminaries on Metric Geometry
@@
where the supremum is taken over all possible finite partitions $0=t_{0}\leq\cdots\leq t_{I}=1$. A curve $\gamma$ is called a geodesic if it satisfies $d(\gamma(s),\gamma(t))=|t-s|d(\gamma(0),\gamma(1))$ for all $s,t\in[0,1]$. The metric space $\mathcal{M}$ is called a geodesic space if, for any two points $\alpha,\beta\in\mathcal{M}$, there exists a geodesic connecting them, denoted $\gamma_{\alpha,\beta}$. If this geodesic is unique for every pair of points in $\mathcal{M}$, the space is referred to as a unique geodesic space (Bridson and Haefliger, 1999).
@@
Example 1 (Networks).
Consider the space of simple, undirected, weighted networks with $m$ nodes. Each network can be uniquely represented by its graph Laplacian matrix. The space of graph Laplacians $\mathcal{L}$ serves as a natural framework for characterizing network structures (Kolaczyk and Csárdi, 2014; Severn et al., 2022; Zhou and Müller, 2022). When endowed with the Frobenius metric, defined as $d_{F}(L_{1},L_{2})=\|L_{1}-L_{2}\|_{F}$ for $L_{1},L_{2}\in\mathcal{L}$, the space $\mathcal{L}$ becomes a unique geodesic space. In this setting, the geodesic between $L_{1}$ and $L_{2}$ is given by the linear interpolation $\gamma_{L_{1},L_{2}}(t)=(1-t)L_{1}+tL_{2}$ for $t\in[0,1]$.
Example 2 (Compositional Data).
@@
Adopting the square-root transformation provides a map from $\Delta^{d-1}$ to the positive orthant of the unit sphere $\mathcal{S}_{+}^{d-1}$ (Scealy and Welsh, 2011), where the induced geodesic distance on $\mathcal{S}_{+}^{d-1}$ is the arc-length distance
@@
with $\mathbf{z}_{1},\mathbf{z}_{2}\in\mathcal{S}_{+}^{d-1}$. Let $\theta=\arccos(\mathbf{z}_{1}^{\prime}\mathbf{z}_{2})$ be the angle between $\mathbf{z}_{1}$ and $\mathbf{z}_{2}$ and $\|\cdot\|$ be the standard Euclidean metric. The geodesic connecting $\mathbf{z}_{1}$ and $\mathbf{z}_{2}$ is then
@@
Example 3 (Symmetric Positive-Definite Matrices).
The space of $m\times m$ symmetric positive-definite (SPD) matrices, denoted by $\mathrm{Sym}^{+}_{m}$, plays a crucial role in various fields such as computer vision (Cherian and Sra, 2016), signal processing (Arnaudon et al., 2013) and neuroimaging (Dryden et al., 2009). Various metrics have been introduced to endow $\mathrm{Sym}^{+}_{m}$ with a meaningful geometric structure, depending on the application. The Frobenius metric $d_{F}$ induces unique geodesics that are simple linear interpolations, $\gamma_{A,B}(t)=(1-t)A+tB$ for $A,B\in\mathrm{Sym}^{+}_{m}$ and $t\in[0,1]$. The space $\mathrm{Sym}^{+}_{m}$ also forms a unique geodesic space under various other metrics such as the Log-Euclidean metric (Arsigny et al., 2007), the power metric family (Dryden et al., 2009) or the Log-Cholesky metric (Lin, 2019). For instance, when adopting the log-Euclidean metric, the unique geodesic connecting $A,B\in\mathrm{Sym}^{+}_{m}$ is given by
@@
Example 4 (One-Dimensional Probability Distributions).
@@
$d_{\mathcal{W}}(\mu,\nu)=\left(\int_{0}^{1}\{F_{\mu}^{-1}(s)-F_{\nu}^{-1}(s)\}^{2}ds\right)^{1/2},$
where $F_{\mu}^{-1}$ and $F_{\nu}^{-1}$ are the quantile functions of the probability measures $\mu$ and $\nu$, respectively. The Wasserstein space $\mathcal{W}$ is a complete, separable and unique geodesic space (Ambrosio et al., 2008), with geodesics given by McCann’s interpolant (McCann, 1997)
@@
Example 5 (Functional Data).
In functional data analysis, one considers data consisting of functions that are considered to be the realizations of a square integrable stochastic process, usually assumed to be defined on a finite interval. Such data are common in the social, physical and life sciences (Ramsay and Silverman, 2005; Hsing and Eubank, 2015; Wang et al., 2016). Typically, functional data are assumed to be situated in the Hilbert space $L^{2}(\mathcal{T})$ of square integrable functions, where $\mathcal{T}$ is a compact interval, most commonly a time domain. In contrast to the other metric spaces we consider in this paper, this space is a vector space and thus permits algebraic operations. It is equipped with the standard $L^{2}$ metric
$d_{L^{2}}(g_{1},g_{2})=\left(\int_{\mathcal{T}}\{g_{1}(t)-g_{2}(t)\}^{2}dt\right)^{1/2},\quad g_{1},g_{2}\in L^{2}(\mathcal{T}),$
@@
3 Geodesic Synthetic Control Methods
3.1 Construction of the estimator
@@
We begin with the conventional synthetic control method (SCM) for Euclidean outcomes. Suppose outcomes $Y_{j,t}\in\mathbb{R}^{q}$ are observed for $J+1$ units over $T$ time periods, indexed by $j=1,\dots,J+1$ and $t=1,\dots,T$. Let $T_{0}>1$ denote the last pre-treatment period, and assume that only unit $j=1$ receives the intervention for $t>T_{0}$. For unit $j$ at time $t$, we denote potential outcomes under treatment and no treatment by $Y_{j,t}(I)$ and $Y_{j,t}(N)$, respectively. We observe $Y_{1,t}=Y_{1,t}(I)$ for $t>T_{0}$ and $Y_{j,t}=Y_{j,t}(N)$ otherwise. The causal effect of interest is
$\tau_{t}=Y_{1,t}(I)-Y_{1,t}(N),\quad t=T_{0}+1,\dots,T.$
In Euclidean settings, the synthetic control estimator constructs a convex combination of the control units that best reproduces the pre-treatment trajectory of the treated unit. Let $\Delta^{J-1}=\{\bm{w}=(w_{2},\ldots,w_{J+1})^{\prime}\in\mathbb{R}^{J}:w_{j}\geq 0,\sum_{j=2}^{J+1}w_{j}=1\}$ denote the probability simplex, and $\overline{\bm{w}}=(\overline{w}_{2},\dots,\overline{w}_{J+1})^{\prime}\in\Delta^{J-1}$ be the optimal weight vector. The synthetic control outcome for the treated unit is defined as
$Y_{1,t}^{(\mathrm{SC})}=\sum_{j=2}^{J+1}\overline{w}_{j}Y_{j,t},\quad t=T_{0}+1,\dots,T,$
where the weights are chosen to minimize the mean squared pre-treatment discrepancy,
$\overline{\bm{w}}\in\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J-1}}\frac{1}{T_{0}}\sum_{t=1}^{T_{0}}\big\|Y_{1,t}-Y_{2:J+1,t}^{(\bm{w})}\big\|^{2},\quad Y_{2:J+1,t}^{(\bm{w})}=\sum_{j=2}^{J+1}w_{j}Y_{j,t}.$
The SCM estimator of the causal effect is then given by
$\tau_{t}^{(\mathrm{SC})}=Y_{1,t}-Y_{1,t}^{(\mathrm{SC})},\quad t=T_{0}+1,\dots,T.$
The construction of SCM relies on two key components: the weighted mean of the control outcomes $Y_{2:J+1,t}^{(\bm{w})}$ and the difference $Y_{1,t}-Y_{1,t}^{(\mathrm{SC})}$. Both are natural in Euclidean space but are not directly available for non-Euclidean outcomes, such as distributions or networks, where algebraic operations like addition, subtraction, or scalar multiplication are not defined. For metric spaces, the appropriate analogue of the weighted mean is the weighted Fréchet mean (Fréchet, 1948; Ginestet et al., 2012), defined for objects $Y_{2,t},\dots,Y_{J+1,t}\in\mathcal{M}$ in a metric space $(\mathcal{M},d)$ as
$Y_{2:J+1,t}^{(\bm{w})}=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,Y_{j,t}),$
which reduces to the weighted Euclidean average when $\mathcal{M}=\mathbb{R}^{q}$ and $d$ is the Euclidean distance. For defining causal effects, differences are also not available outside Euclidean space. However, when the outcome space is geodesic, a suitable generalization exists: in Euclidean space, the difference $\beta-\alpha$ can be equivalently interpreted as movement along the straight line (geodesic) from $\alpha$ to $\beta$, which motivates defining causal effects via geodesics in geodesic metric spaces (Kurisu et al., 2024).
Accordingly, in a unique geodesic space $(\mathcal{M},d)$ we define the causal effect of the intervention as the geodesic connecting the untreated and treated potential outcomes,
$\tau_{t}=\gamma_{Y_{1,t}(N),Y_{1,t}(I)},\quad t=T_{0}+1,\dots,T.$ (1)
The uniqueness of geodesics ensures that $\tau_{t}$ is well defined. This definition preserves the interpretation of $\tau_{t}$ as encoding both the distance and direction of change.
To estimate $\tau_{t}$, we extend the weight construction from SCM using weighted Fréchet means. The synthetic control outcome for the treated unit is
$Y_{1,t}^{(\mathrm{GSC})}=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}\overline{w}_{j}d^{2}(\nu,Y_{j,t}),$
with weights chosen as
@@
$\overline{\bm{w}}\in\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J-1}}\frac{1}{T_{0}}\sum_{t=1}^{T_{0}}d^{2}\!\left(Y_{1,t},Y_{2:J+1,t}^{(\bm{w})}\right),\quad Y_{2:J+1,t}^{(\bm{w})}=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,Y_{j,t}).$ (2)
The geodesic synthetic control (GSC) estimator of the treatment effect is then